# Multiple Information Prompt Learning for Cloth-Changing Person Re-Identification

### v1

@@ III-D Loss Function
Where $p^{img}$ denotes output class probabilities of the backbone, $p^{bio}$ denotes output class probabilities of the BGA, Kullback-Leibler (KL) divergence [41] is used to quantify the matching degree of the two output probabilities, and $\mathcal{D}_{KL}$ represents the KL distance. In this way, the backbone is endowed with the ability to extract the biological knowledge with strong identity correlation from the features, so as to improve the feature robustness of person reID in the cloth-changing scenario.
@@ IV Experiments And Discussion
TABLE I: Performance evaluation and comparison on five public cloth-changing datasets, where the bold values indicate the best performance in each column, and underlined values indicate the optimal performance of the existing methods.
@@
To evaluate the performance of the MIPL method, we performed experiments using five public cloth-changing person ReID datasets: PRCC [34], LTCC [24], Celeb-reID [16], Celeb-reID-light [15], and CSCC[33]. The remainder of this section is organized as follows: 1) the competing methods used in our experiments are listed, 2) the implementation details are described, and 3) the performance evaluations and comparisons based on these five public datasets are described.
@@ IV-B Competitors
The task of CC-ReID is a new and challenging topic that has also aroused the interest of researchers in related fields in recent years. In our experiments, the latest and popular references were utilized as our competitors, including FSAM (CVPR 2021) [14], 3DSL (CVPR 2021) [1], MAC-DIM (TMM 2021) [2], LaST (TCSVT 2021) [27], GI-ReID (CVPR 2022) [18], CAL (CVPR 2022) [8], MVSE (ACM MM 2022) [7], UCAD (IJCAI 2022) [33], Pos-Neg (TIP 2022) [22], DCR-ReID (TCSVT 2023) [3], AIM (CVPR 2023) [35], CCFA (CVPR 2023) [10], SAVS (TNNLS 2023) [6], IMS+GEP (TMM 2023) [42], PGAL (TMM 2023) [23], CT-Net (TMM 2023) [32], ACID (TIP 2023) [36], SCNet (ACM MM 2023) [9], CSCL (ACM MM 2023) [30], DLAW (TIP 2023) [22], and MBUNet (TIP 2023) [39]. Additionally, in the CC-ReID task, traditional person ReID algorithms, such as ResNet50 (CVPR 2016) [12], ViT-B/16 (ICLR 2021) [4], PCB (ECCV 2018) [28], and MGN (ACM MM 2018) [29], are often employed. In our experiments, we also compared MIPL with them. More information about these competitors can be obtained in the related work section.

### v2

@@ III-B Bio-guided Attention (BGA)
Where $p^{img}$ denotes output class probabilities of the backbone, $p^{bio}$ denotes output class probabilities of the BGA, Kullback-Leibler (KL) divergence [46] is used to quantify the matching degree of the two output probabilities, and $\mathcal{D}_{KL}$ represents the KL distance.
In this way, the backbone is endowed with the ability to extract the biological knowledge with strong identity correlation from the features, so as to improve the feature robustness of person reID in the cloth-changing scenario.
III-C Dual-length Hybrid Patch (DHP)
III-D Loss Function
$\mathcal{L}_{Stage1}=\mathcal{L}_{i2t}^{ID}(y_{i})+\mathcal{L}_{i2t}^{C}(y_{c})+\mathcal{L}_{t2i}^{ID}(y_{i})+\mathcal{L}_{t2i}^{C}(y_{c})$ (10)
which includes the image-text contrastive loss $\mathcal{L}_{{i2t}}$ and the text-image contrastive loss $\mathcal{L}_{{t2i}}$, where $ID$ and $C$ represent identity and clothing dependency respectively, and $y_{i}$ represents identity labels, and $y_{c}$ represents clothing labels.
The second training stage. In this stage, both the text prompt and the text encoder are frozen and only the image encoder is optimized, we follow previous ReID work [15, 22, 49] to calculate cross-entropy loss and triplet loss to optimize the image encoder. For the cross-entropy loss $\mathcal{L}_{ce}$ and the triplet loss $\mathcal{L}_{tri}$, they are calculated as follows:
$\mathcal{L}_{ce}=\frac{1}{B}\sum_{i=1}^{B}-\log p(x\mid y)$ (11)
$\mathcal{L}_{tri}=\frac{1}{B}\sum_{i=1}^{B}\max\{m+d_{p}-d_{n},0\}$ (12)
Where $p(x\mid y)$ is the predicted probability that the sample $x$ belongs to the ground truth $y$. $d_{p}$ and $d_{n}$ are the feature distances of positive pair and negative pair, and $m$ is the margin of $\mathcal{L}_{tri}$.
In addition, in order to strip the clothing interference information, we design the clothing stripping loss to guide the model to decouple and strip the clothing information. In order to effectively use biological information to enhance features, the biological guided loss is used to prompt the model to learn biological knowledge with a strong identity correlation. Thus, the loss function in the second stage is defined by
@@ IV Experiments And Discussion
TABLE I: Performance evaluation and comparison on five public cloth-changing datasets, where the bold values indicate the best performance in each column, and underlined values indicate the optimal performance of the existing methods. The MIPL* denotes the scheme of overlapping patches is adopted.
@@
To evaluate the performance of the MIPL method, we performed experiments using five public cloth-changing person ReID datasets: PRCC [38], LTCC [27], Celeb-reID [18], Celeb-reID-light [17], and CSCC[37]. The remainder of this section is organized as follows: 1) five public cloth-changing person ReID datasets are introduced, 2) the competing methods used in our experiments are listed, 3) the implementation details are described, 4) the performance evaluations and comparisons based on these five public datasets are described, and 5) convergence analysis.
@@ IV-B Competitors
The task of CC-ReID is a new and challenging topic that has also aroused the interest of researchers in related fields in recent years. In our experiments, the latest and popular references were utilized as our competitors, including FSAM (CVPR 2021) [16], 3DSL (CVPR 2021) [2], MAC-DIM (TMM 2021) [3], LaST (TCSVT 2021) [31], GI-ReID (CVPR 2022) [20], CAL (CVPR 2022) [10], MVSE (ACM MM 2022) [9], UCAD (IJCAI 2022) [37], Pos-Neg (TIP 2022) [19], DCR-ReID (TCSVT 2023) [4], AIM (CVPR 2023) [40], CCFA (CVPR 2023) [12], SAVS (TNNLS 2023) [8], IMS+GEP (TMM 2023) [47], PGAL (TMM 2023) [25], CT-Net (TMM 2023) [36], ACID (TIP 2023) [41], SCNet (ACM MM 2023) [11], CSCL (ACM MM 2023) [34], DLAW (TIP 2023) [24], and MBUNet (TIP 2023) [44]. Additionally, in the CC-ReID task, traditional person ReID algorithms, such as ResNet50 (CVPR 2016) [14], ViT-B/16 (ICLR 2021) [6], PCB (ECCV 2018) [32], and MGN (ACM MM 2018) [33], are often employed. In our experiments, we also compared MIPL with them. In addition, the introduction of non-RGB modal information can provide richer information to the model. For example, 3DSL, FSAM, MAC-DIM, GI-ReID, DCR-ReID, and PGAL all use different modal information. It is worth noting that MIPL is the first method to introduce text information as a prompt into the task of CC-ReID. More information about these competitors can be obtained in the related work section.